OpenStudy (anonymous):

How many solutions are there to the quadratic equation 4x^2 + 8x + 3 = 0? 0 1 2 none of the above

jimthompson5910 (jim_thompson5910):

Do you know the discriminant formula?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

If you have ax^2 + bx + c = 0 The discriminant formula is D = b^2 - 4ac

jimthompson5910 (jim_thompson5910):

If D < 0, then you'll have no solutions If D = 0, then you'll have one solution If D > 0, then you'll have two solutions

jimthompson5910 (jim_thompson5910):

Does that help?

OpenStudy (anonymous):

im confused

jimthompson5910 (jim_thompson5910):

Here's a similar example

OpenStudy (amistre64):

all quadratics have "2" solutions ....

OpenStudy (anonymous):

maybe she should try to factorize,then she'll get it

jimthompson5910 (jim_thompson5910):

Problem: Find the discriminant of 3x^2 + 11x - 4 = 0 ------------------------------------------- Steps: 3x^2 + 11x - 4 = 0 is in the form ax^2 + bx + c = 0 where a = 3 b = 11 c = -4 Plug all this into the discriminant formula D = b^2 - 4ac to get D = b^2 - 4ac D = 11^2 - 4(3)(-4) D = 121 + 48 D = 169 Since D > 0, this means that there are 2 real solutions -------------------------------------------- Answer: The equation 3x^2 + 11x - 4 = 0 has two real solutions. These solutions are also distinct.

jimthompson5910 (jim_thompson5910):

Keep in mind that the answer above is the answer to the example problem and not your specific problem (but they're similar)